# Does provable absence of barren plateaus imply classical simulability? Or, why we need to rethink variational quantum computing

### v2

@@ Abstract
A large amount of effort has recently been put into understanding the barren plateau phenomenon. In this perspective article, we face the increasingly loud elephant in the room and ask a question that has been hinted at by many but not explicitly addressed: Can the structure that allows one to avoid barren plateaus also be leveraged to efficiently simulate the loss classically? We present strong evidence that commonly used models with provable absence of barren plateaus are also classically simulable, provided that one can collect some classical data from quantum devices during an initial data acquisition phase. This follows from the observation that barren plateaus result from a curse of dimensionality, and that current approaches for solving them end up encoding the problem into some small, classically simulable, subspaces. Thus, while stressing quantum computers can be essential for collecting data, our analysis sheds serious doubt on the non-classicality of the information processing capabilities of parametrized quantum circuits for barren plateau-free landscapes. We end by discussing caveats in our arguments, the role of smart initializations and the possibility of provably superpolynomial, or simply practical, advantages from running parametrized quantum circuits.
@@ I Introduction
In recent years, the initial excitement attracted by variational quantum algorithms [1, 2, 3] and quantum machine learning [4, 5, 6, 7, 8, 9] has been tempered by the barren plateau phenomenon [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Namely, there is a growing awareness that a large class of quantum learning architectures exhibit loss function landscapes that concentrate exponentially in system size towards their mean value. On such landscapes, exponential resources are required for training, prohibiting the successful scaling of variational quantum algorithms. Hence, identifying architectures and training strategies that provably do not lead to barren plateaus has become a highly active area of research. Examples of such strategies include shallow circuits with local measurements [22, 13, 23, 24, 25, 26], dynamics with small Lie algebras [20, 48, 49, 50, 51, 52, 56], identity initializations [37, 53, 58], embedding symmetries into the circuit’s architecture [59, 60, 61, 62, 63, 64], and certain classes of quantum generative models [31, 30, 27].
@@
However, these strategies all, in some sense, make use of some simple underlying structure of the problem. This provokes the question: Could the very same structure that allows one to provably avoid barren plateaus be leveraged to efficiently simulate the loss function classically? Here we argue that the answer to this question is ‘yes’. Specifically, we claim that loss landscapes which provably do not exhibit barren plateaus can be simulated using a classical algorithm that runs in polynomial time. Importantly, this simulation might still necessitate the use of a quantum computer during an initial data acquisition phase [65, 66, 67, 68], but it does not require parametrized quantum circuits implemented on a quantum device nor hybrid quantum-classical optimization loops. These arguments can be understood as a form of dequantization of the information processing capabilities of variational quantum circuits in barren plateau-free landscapes.
@@
Core to our argument is the observation that any loss based on evolving an initial state $\rho$ through a parameterized quantum circuit $U(\bm{\theta})$ and then estimating the expectation value of an observable $O$ can be written as an inner product between the Heisenberg evolved observable $U(\bm{\theta})^{\dagger}OU(\bm{\theta})$ and the state $\rho$. Given that both of these objects live in the exponentially large vector space of operators, one can generally expect this overlap to be –on average over $\bm{\theta}$– exponentially small. This is the essence of the barren plateaus phenomenon—the curse of dimensionality. If, however, the evolved observable is confined to a polynomially large subspace, then the loss becomes the inner product between two objects in this reduced space and can therefore avoid barren plateaus. But, in this case one can also simulate the loss by representing the initial state, circuit, and measurement operator as polynomially large objects contained in, and acting on, the small subspace.
@@
Our general argument is supported by an analysis of widely used schemes through which we show that all considered methods for avoiding barren plateaus can be efficiently classically simulated. Fundamentally, it is the very proof of absence of barren plateaus that allows us to identify the polynomially-sized subspaces in which the relevant part of the computation lives. Using this information, we can then determine the set of expectation values one needs to estimate (either classically or quantumly) to enable classical simulations. In the latter case the quantum computer is still essential, but is accessed non-adaptively to generate a classical surrogate rather than via a hybrid-quantum classical optimization loop.
@@
Secondly, while for all our case studies it is possible to identify the ingredients necessary for simulation, we do not prove that this will always be possible. Thus, in principle, there could be models for which the landscape is free of barren plateaus, and yet we do not know how to simulate it. This could arise for sub-regions of a landscape which could be explored via smart initialization strategies, when the small subspace is otherwise unknown, or even when the problem lives in the full exponential space but is highly structured. Indeed, we provide an explicit (but highly contrived) construction for the latter.
@@ II Definitions for barren plateaus and simulability
Here, $\rho\in\mathcal{B}$ is an $n$-qubit input state belonging to the set of bounded operators $\mathcal{B}$ acting on a $2^{n}$-dimensional Hilbert space $\mathcal{H}$, $O\in\mathcal{B}$ is some Hermitian operator (with $\norm{O}_{\infty}\leqslant 1$), $U(\bm{\theta})$ is the parametrized quantum circuit, and $\bm{\theta}$ is a set of trainable parameters. Losses of this form can be used to tackle a wide range of problems through different choices of $\rho$, $O$ and $U(\bm{\theta})$. We note that while algorithms can employ more general loss functions that require computing multiple such quantities (e.g., by sending different states through the circuit, or by estimating the expectation value of several operators), we will focus on the fundamental case where the loss is given by Eq. (1), as the lessons derived here can be extrapolated to other scenarios.
Figure 1: Schematic description of the simulation classes. a) Problems in ${\rm CSIM}$ are those for which there exists a fully classical algorithm that takes as input the problem description and estimates the loss function in polynomial time with a classical computer. Here, access to a quantum computer is not needed. b) Problems in ${\rm CSIM}_{\mathrm{QE}}$ are those where one is allowed access to a quantum computer for an initial data acquisition phase that takes no more than polynomial time. At the end of this phase, access to the quantum computer is no longer allowed. A classical algorithm then takes the problem’s description, and the data obtained from the quantum device, to estimate the loss function in polynomial time. Note that one can compute problems in ${\rm CSIM}_{\mathrm{QE}}$ without needing to run a parametrized quantum circuit on the quantum hardware. c) Problems in ${\rm QSIM}$ are those where one allows ‘on demand’ access to a quantum computer. Here, one usually implements the parametrized quantum circuit on the device.
@@
Over the past few years, there has been a tremendous amount of work put forward to understand if the loss functions in a given problem class are trainable. Several sources of untrainability have been detected [7] such as the presence of sub-optimal local minima [70, 71, 72, 73, 20] and expressivity limitations [74]. However, the vast majority of trainability analysis has been concentrated around the barren plateau phenomenon [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 30, 27, 28, 29, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. When a problem exhibits a barren plateau, its loss function becomes—on average—exponentially concentrated with the system size [75, 22]. We say a class $\mathcal{C}$ of problems is provably barren plateau-free, i.e., $\mathcal{C}\in\overline{{\rm BP}}$, if one can show that
@@
To better tackle this question, we then first need to define what it means to compute a loss function, and also what we understand by classical simulability. In particular, since the notion of barren plateaus is an average statement over the landscape, in this context the natural notions for computing and simulating the loss will also be average ones. However, stronger notions of simulability, where one guarantees that the loss can be computed for all points on the landscape, will also be discussed below.
First, let us define the task of computing a loss function, such as that in Eq. (1), for a given problem. We will say that an algorithm can compute the loss functions in an instance $\mathcal{I}$ if, with high probability for $\bm{\theta}\sim\mathcal{P}$, it can implement a function ~$\tilde{\ell}_{\bm{\theta}}(\rho,O)$ approximating the loss up to error $\epsilon$, i.e.,
(~$\left(\tilde{\ell}_{\bm{\theta}}(\rho,O)-\ell_{\bm{\theta}}(\rho,O)\right)^{2}\leqslant\epsilon\,.$ (3)
@@
As previously mentioned, the above definition guarantees that we can compute the loss function with high probability when the parameters are sampled according to $\mathcal{P}$. One might also be interested in computing the loss given any parameter settings. Thus, we will also consider a stronger version where an algorithm can compute the loss function in an instance $\mathcal{I}$ if for all $\bm{\theta}$, it can implement a function ~$\tilde{\ell}_{\bm{\theta}}(\rho,O)$ approximating the loss up to error $\epsilon$.
@@
We also consider classical simulation algorithms enhanced by polynomial-size data obtained from quantum experiments, which we denote as Classical Simulation enhanced with Quantum Experiments (${\rm CSIM}_{\mathrm{QE}}$). In this case, one is given a problem instance and is allowed to use a quantum computer for an initial data acquisition phase, which takes no more than polynomial time. During this phase, one can prepare copies of the initial quantum state, apply some operations, and independently measure them via some efficient tomographic or classical shadow techniques [78, 66, 79]. Such a procedure can be used to obtain an efficiently storable classical representation (i.e., storable in string of size in $\mathcal{O}(\operatorname{poly}(n))$) of the state, unitary or measurement operator. Once this initial phase is over, one cannot access the quantum device anymore, and the computation of $\ell_{\bm{\theta}}(\rho,O)$ has to be done purely classically. A problem $\mathcal{C}$ is in ${\rm CSIM}_{\mathrm{QE}}$ if a polynomial-time classical algorithm, which can utilize data obtained from quantum devices in an initial data acquisition phase, can compute every instance in $\mathcal{C}$ [^1] Note that the problem of estimating a loss with data from a quantum computer can be cast into a decision problem. In this case, ${\rm CSIM}$ can be related to BPP, ${\rm CSIM}_{\mathrm{QE}}$ is closely connected to BPP/Samp [65] and even more closely connected to BPP/qgenpoly [80], and ${\rm QSIM}$ is related to BQP. Speaking more informally, algorithms in ${\rm CSIM}_{\mathrm{QE}}$ are also often described as classical surrogate methods..
We stress that all problems in ${\rm CSIM}$ and ${\rm CSIM}_{\mathrm{QE}}$ do not require running a parametrized quantum circuit on the quantum computer. Quantum resources are either entirely unnecessary (${\rm CSIM}$), or used only in some initial data acquisition phase (${\rm CSIM}_{\mathrm{QE}}$).
Figure 2: Class inclusions. We show the inclusions between the classes ${\rm CSIM}$, ${\rm CSIM}_{\rm QE}$ and ${\rm QSIM}$. The region outside of ${\rm CSIM}$ (highlighted with stripes) corresponds to problems where a quantum advantage could potentially be achieved.
@@
${\rm CSIM}\subseteq{\rm CSIM}_{\mathrm{QE}}\subseteq{\rm QSIM}\,.$ (4)
For instance, ${\rm CSIM}\subseteq{\rm CSIM}_{\mathrm{QE}}$ because if the loss is fully classically simulable then it can be estimated by simply skipping the quantum computer. The rest of the inclusions follow similarly. Clearly, QSIM is not the largest possible set, and any loss that requires exponential time to estimate, even with a quantum computer, is beyond QSIM.
A quantum advantage is possible for problems where the loss can be simulated only if we have access to a quantum computer, i.e., for any problem in ${\rm QSIM}\cap\overline{{\rm CSIM}}\,$. In fact, any problem where the loss is in ${\rm CSIM}_{\mathrm{QE}}$ but not ${\rm CSIM}$ is already capable of a quantum advantage as it requires a quantum device. The problems that fall in ${\rm CSIM}_{\mathrm{QE}}\cap\overline{{\rm CSIM}}$ may well be the most suited for implementation on a near-term quantum computer as the data acquisition phase could be less noisy than fully implementing the parametrized quantum circuit [81, 82, 83, 65, 80, 84].
@@
$\overline{{\rm BP}}\overset{?}{\subset}{\rm CSIM}\,\,\mathrm{or}\,\,\overline{{\rm BP}}\overset{?}{\subset}{\rm CSIM}_{\mathrm{QE}}\,.$ (5)

### v3

@@ Abstract
A large amount of effort has recently been put into understanding the barren plateau phenomenon. In this perspective article, we face the increasingly loud elephant in the room and ask a question that has been hinted at by many but not explicitly addressed: Can the structure that allows one to avoid barren plateaus also be leveraged to efficiently simulate the loss classically? We collect evidence—on a case-by-case basis—that many commonly used models whose loss landscapes avoid barren plateaus can also admit classical simulation, provided that one can collect some classical data from quantum devices during an initial data acquisition phase. This follows from the observation that barren plateaus result from a curse of dimensionality, and that current approaches for solving them end up encoding the problem into some small, classically simulable, subspaces. Thus, while stressing that quantum computers can be essential for collecting data, our analysis sheds doubt on the information processing capabilities of many parametrized quantum circuits with provably barren plateau-free landscapes. We end by discussing the (many) caveats in our arguments including the limitations of average case arguments, the role of smart initializations, models that fall outside our assumptions, the potential for provably superpolynomial advantages and the possibility that, once larger devices become available, parametrized quantum circuits could heuristically outperform our analytic expectations.
@@ I Introduction
In recent years, the initial excitement attracted by variational quantum algorithms [1, 2, 3] and quantum machine learning [4, 5, 6, 7, 8, 9] has been tempered by the barren plateau phenomenon [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Namely, there is a growing awareness that a large class of quantum learning architectures exhibit loss function landscapes that concentrate exponentially in system size towards their mean value. On such landscapes, exponential resources are required for training, prohibiting the successful scaling of variational quantum algorithms. Hence, identifying architectures and training strategies that provably do not lead to barren plateaus has become a highly active area of research. Examples of such strategies include shallow circuits with local measurements [24, 15, 25, 26, 27, 28], dynamics with small Lie algebras [22, 50, 51, 52, 53, 54, 58, 60], identity initializations [39, 55, 61], embedding symmetries into the circuit’s architecture [62, 63, 64, 65, 66, 67, 68, 69], adding non-unital noise or intermediate measurements [56, 70, 71, 72, 73], and certain classes of quantum generative models [33, 32, 29].
@@
However, these strategies all, in some sense, make use of some simple underlying structure of the problem. This provokes the question: Could the very same structure that allows one to provably avoid barren plateaus be leveraged to efficiently simulate the loss function classically? Here we argue that the answer to this question is “Yes and No”. Specifically, we argue and present strong evidence that a wide class of loss landscapes which provably do not exhibit barren plateaus can be simulated using either a classical algorithm or what we call a “quantum-enhanced” classical algorithm that runs in polynomial time. In the latter case, this simulation still necessitates the use of a quantum computer during an initial data acquisition phase [74, 75, 76, 77], but it does not require hybrid quantum-classical optimization loops. These arguments can be understood as a soft form of dequantization of the information processing capabilities of many variational quantum circuits in barren plateau-free landscapes.
@@
Core to our argument is the observation that any loss based on evolving an initial state $\rho$ through a parameterized quantum circuit $U(\bm{\theta})$ and then estimating the expectation value of an non-trivial observable $O$ can be written as an inner product between the Heisenberg evolved observable $U(\bm{\theta})^{\dagger}OU(\bm{\theta})$ and the state $\rho$. Given that both of these objects live in the exponentially large vector space of operators, one can generally expect this overlap to be –on average over $\bm{\theta}$– exponentially small. This is the essence of the barren plateaus phenomenon—the curse of dimensionality. If, however, the evolved observable is confined to a polynomially large subspace, then the loss becomes the inner product between two objects in this reduced space and can therefore avoid barren plateaus. But, in this case one can also simulate the loss by representing the initial state, circuit, and measurement operator as polynomially large objects contained in, and acting on, the small subspace.
@@
Our general argument is supported by an analysis of widely used schemes through which we show that all considered methods for avoiding barren plateaus can be efficiently classically simulated, as well as by recent works which explicitly perform the classical simulations [78, 79, 80, 81, 70, 82, 83, 84]. Fundamentally, it is the very proof of absence of barren plateaus that allows us to identify the polynomially-sized subspaces in which the relevant part of the computation lives. Using this information, we can then determine the set of expectation values one needs to estimate (either classically or quantumly) to enable classical simulations. In the latter case the quantum computer is still essential, but is accessed non-adaptively to generate a classical surrogate rather than via a hybrid optimization loop.
@@
Secondly, while for all our case studies it is possible to identify the ingredients necessary for simulation, we do not prove that this will always be possible. Thus, in principle, there could be models for which the landscape is free of barren plateaus, and yet we do not know how to simulate it. This could arise for sub-regions of a landscape which could be explored via smart initialization strategies, when the small subspace is otherwise unknown, or even when the problem lives in the full exponential space but is highly structured. Indeed, we provide an explicit (but highly contrived) construction for the latter and raise the possibility that more such cases might be found heuristically once larger quantum devices become available.
@@ II Definitions for barren plateaus and simulability
Here, $\rho\in\mathcal{B}$ is an $n$-qubit input state belonging to the set of bounded operators $\mathcal{B}$ acting on a $2^{n}$-dimensional Hilbert space $\mathcal{H}$, $O\in\mathcal{B}$ is some non-trivial Hermitian operator (with $\norm{O}_{\infty}\leqslant 1$), $U(\bm{\theta})$ is the parametrized quantum circuit, and $\bm{\theta}$ is a set of trainable parameters. Losses of this form can be used to tackle a wide range of problems through different choices of $\rho$, $O$ and $U(\bm{\theta})$. We note that while algorithms can employ more general loss functions that require computing multiple such quantities (e.g., by sending different states through the circuit, or by estimating the expectation value of several operators), we will focus on the fundamental case where the loss is given by Eq. (1), as the lessons derived here can be extrapolated to other scenarios.
Figure 1: Schematic description of the simulation classes. a) Problems in $\rm CSIM$} are those for which there exists a fully classical algorithm that takes as input the problem description and estimates the loss function in polynomial time with a classical computer. Here, access to a quantum computer is not needed. b) Problems in $\rm QESIM$} are those where one is allowed access to a quantum computer for an initial data acquisition phase that takes no more than polynomial time. At the end of this phase, access to the quantum computer is no longer allowed. A classical algorithm then takes the problem’s description, and the data obtained from the quantum device, to estimate the loss function in polynomial time. Note that one can compute problems in $\rm QESIM$} without needing to run a parametrized quantum circuit on the quantum hardware. c) Problems in $\rm QSIM$} are those where one allows ‘on demand’ access to a quantum computer. Here, one usually implements the parametrized quantum circuit on the device.
@@
Over the past few years, there has been a tremendous amount of work put forward to understand if the loss functions in a given problem class based on variational circuits are trainable (and we refer the reader to Ref. [86] for a subtle discussion of what trainable, or even variational, even means). Several sources of untrainability have been detected [7] such as the presence of sub-optimal local minima [87, 88, 89, 90, 22] and expressivity limitations [91]. However, the vast majority of trainability analysis has been concentrated around the barren plateau phenomenon [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 32, 29, 30, 31, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59] (see Ref. [11] for a review on barren plateaus). When a problem exhibits a barren plateau, its loss function becomes—on average—exponentially concentrated with the system size [10, 24]. We say a class $\mathcal{C}$ of problems is provably barren plateau-free, i.e., $\mathcal{C}\in\overline{{\rm BP}}$, if one can show that
@@
To better tackle this question, we then first need to define what it means to compute a loss function, and also what we understand by classical simulability. In particular, since the notion of barren plateaus is an average statement over the landscape, in this context the natural notions for computing and simulating the loss will also be average ones. However, stronger notions of simulability, such as one where one guarantees that the loss can be computed for all points on the landscape, will also be discussed below.
First, let us define the task of computing a loss function, such as that in Eq. (1), for a given problem. We will say that an algorithm can compute the loss functions in an instance $\mathcal{I}$ if, with high probability $\bm{\theta}\sim\mathcal{P}$, it can implement a function $\tilde{\ell}_{\bm{\theta}}(\rho,O)$ approximating the loss up to error $\epsilon$, i.e.,
$\left(\tilde{\ell}_{\bm{\theta}}(\rho,O)-\ell_{\bm{\theta}}(\rho,O)\right)^{2}\leqslant\epsilon\,.$ (3)
@@
As previously mentioned, the above definition guarantees that we can compute the loss function with high probability when the parameters are sampled according to $\mathcal{P}$. While simulating–with high probability–random point in the landscape is typically easy (see for instance [79]) it may not be particularly useful as it does not necessarily allow us to train the circuit. One might also be interested in computing the loss given any parameter settings. Thus, we will also consider a stronger version where an algorithm can compute the loss function in an instance $\mathcal{I}$ if for all $\bm{\theta}$, it can implement a function $\tilde{\ell}_{\bm{\theta}}(\rho,O)$ approximating the loss up to error $\epsilon$. Of course, more pragmatic definitions exist where we do not care about approximating the loss function in average, or across all points, but where we instead focus on actually solving the optimization task at hand. As such, one may wish to obtain a $\tilde{\ell}_{\bm{\theta}}(\rho,O)$ such that the solution to the optimization problem $\operatorname*{arg\,min}_{\bm{\theta}}\tilde{\ell}_{\bm{\theta}}(\rho,O)$ is as good (according to some metric) as that obtained by solving $\operatorname*{arg\,min}_{\bm{\theta}}\ell_{\bm{\theta}}(\rho,O)$. We will not focus on this case here but we note that its exploration is an open area of research [94, 78].
@@
We also consider classical simulation algorithms enhanced by polynomial-size data obtained from quantum experiments, which we denote as Quantum Enhanced Classical Simulation ($\rm QESIM$}), which morally means that the quantum landscape can be classically surrogated. In this case, one is given a problem instance and is allowed to use a quantum computer for an initial data acquisition phase, which takes no more than polynomial time (and uses no more than polynomial-memory). During this phase, one can prepare copies of the initial quantum state, apply some operations, and independently measure them via some efficient tomographic or classical shadow techniques [95, 75, 96]. Such a procedure can be used to obtain an efficiently storable classical representation (i.e., storable in string of size in $\mathcal{O}(\operatorname{poly}(n))$) of the state, unitary or measurement operator. Once this initial phase is over, one cannot access the quantum device anymore, and the computation of $\ell_{\bm{\theta}}(\rho,O)$ has to be done purely classically. A problem $\mathcal{C}$ is in $\rm QESIM$} if a polynomial-time classical algorithm, which can utilize data obtained from quantum devices in an initial data acquisition phase, can compute every instance in $\mathcal{C}$. Here we note that the problem of estimating a loss with data from a quantum computer can be cast into a decision problem. In this case, $\rm CSIM$} can be related to BPP, $\rm QESIM$} is closely connected to BPP/Samp [74] and even more closely connected to BPP/qgenpoly [97], and $\rm QSIM$} is related to BQP. Speaking more informally, algorithms in $\rm QESIM$} are also often described as classical surrogate methods.
We stress that all problems in $\rm CSIM$} and $\rm QESIM$} do not require running a parametrized quantum circuit on the quantum computer. Quantum resources are either entirely unnecessary ($\rm CSIM$}), or used only in some initial data acquisition phase ($\rm QESIM$}).
Figure 2: Class inclusions. We show the inclusions between the classes $\rm CSIM$}, $\rm QESIM$} and $\rm QSIM$}. The region outside of $\rm CSIM$} (highlighted with stripes) corresponds to problems where a quantum advantage could potentially be achieved.
@@
${\rm CSIM}\subseteq{\rm QESIM}\subseteq{\rm QSIM}\,.$ (4)
For instance, ${\rm CSIM}\subseteq{\rm QESIM}$ because if the loss is fully classically simulable then it can be estimated by simply skipping the quantum computer. The rest of the inclusions follow similarly. Clearly, QSIM is not the largest possible set, and any loss that requires exponential time to estimate, even with a quantum computer, is beyond QSIM.
A quantum advantage is possible for problems where the loss can be simulated only if we have access to a quantum computer, i.e., for any problem in ${\rm QSIM}\cap\overline{{\rm CSIM}}\,$. In fact, any problem where the loss is in $\rm QESIM$} but not $\rm CSIM$} is already capable of a quantum advantage as it requires a quantum device. The problems that fall in ${\rm QESIM}\cap\overline{{\rm CSIM}}$ may well be the most suited for implementation on a near-term quantum computer as the data acquisition phase could be less noisy than fully implementing the parametrized quantum circuit [98, 99, 100, 74, 97, 101].
@@
$\overline{{\rm BP}}\overset{?}{\subset}{\rm CSIM}\,\,\mathrm{or}\,\,\overline{{\rm BP}}\overset{?}{\subset}{\rm QESIM}\,.$ (5)